# Road Graph Generator: Mapping roads at construction sites from GPS data

### v2

@@ Abstract
We propose a new method for inferring roads from GPS trajectories to map construction sites. This task presents a unique challenge due to the erratic and non-standard movement patterns of construction machinery, which significantly diverge from typical vehicular traffic on established roads. Our proposed method first identifies intersections in the road network that serve as critical decision points, and then connects them with edges to produce a graph, which can subsequently be used for planning and task-allocation. We demonstrate the approach by mapping roads at a real-life construction site in Norway. The method is validated on four increasingly complex segments of the map. In our tests, the method achieved perfect accuracy in detecting intersections and inferring roads in data with no or low noise, while its performance was reduced in map areas with significant noise and consistently missing GPS updates.
@@ 1 Motivation
Reducing carbon emissions is a critical goal in mitigating the far-reaching impacts of climate change. In Norway, the building and construction sector contributes directly and indirectly to 15% of total greenhouse gas emissions (2019), with construction vehicles accounting for 1.5% of the total emissions (2021) BAE (1, 2).
In managing construction projects, it is crucial to effectively handle the complex interplay of resources, including personnel, materials, and machinery, to ensure their optimal interaction in terms of type, quantity, location, and timing Dardouri2023 (3). One viable strategy to address this issue is to maximize the utilization of heavy machinery, such as dump trucks and excavators, through optimal coordination and task allocation. According to Skanska Norge AS, the largest construction contractor in Norway, mass transporting units are idling 40–60% of the time during projects. This is a consequence of the conventional planning approach where excavators are considered more expensive than mass transportation units, and hence the latter are excessively deployed to avoid idling excavators, often leading to over-capacity and idling of transportation units instead. Improved coordination on construction sites, e.g., optimized real-time task allocation for dumpers, can reduce this over-capacity, leading to lower greenhouse gas emissions and saving both time and money. This is a critical improvement given the increasing size of construction projects involving hundreds of mass transporting units.
The first step in optimizing these operations is acquiring an up-to-date map of the construction site. It is imperative that the map generation be automated given the constantly evolving nature of the construction site, and because manual mapping would be impractical and deprioritized, as emphasized in interviews with drivers, foremen and construction site coordinators. Herein, we address this need by proposing an algorithm for road inference from Global Positioning System (GPS) trajectories recorded by vehicles on the construction site. Our algorithm generates a graph consisting of road intersections (nodes) and road centerlines (edges), which is a formulation that conveniently allows for the subsequent use of standard resource planning and optimization techniques. Furthermore, utilizing movement data, as opposed to e.g., aerial images facilitates identification of temporary roads which are typically less visible or distinguishable in the terrain and does not require frequent collection of aerial images to capture the dynamic changes.
@@ 2 Related work
Automated road inference plays a pivotal role in various applications, such as urban planning, transportation management, and autonomous driving. The primary machine learning approaches for map generation are either from images, e.g., aerial photos, drone-, or satellite imagery Li2019 (4, 5, 6, 7, 8), or from movement data, such as GPS trajectories 10.1145/3234692 (9, 10, 11). The latter approach becomes particularly advantageous when paths are not easily distinguishable in images, e.g., blend into the surrounding environment. Moreover, GPS data collection is cost-effective compared to aerial or satellite imagery, especially for repeated acquisitions over time, since it can leverage existing devices like mobile phones.
@@
In the context of optimizing vehicle operations at construction sites, it is advantageous to model the road network as a graph. In such model, intersections are treated as key decision points, represented as nodes, while the roads themselves are depicted as the connecting edges between these nodes. Moreover, graph representation is more compact compared to alternatives like raster images. We therefore direct our attention to works that detect road intersections and infer road centerlines, i.e., central paths along the roads.
@@
In CellNet 10.1145/3234692 (9), the approach involves dividing the map into a grid and applying the mean-shifting algorithm on all GPS points within each cell, which results in identifying intersections, if they are present within the cell, or pinpointing the center of the stream of routes otherwise. To filter out the latter, each potential intersection is verified to have at least three outgoing streams of routes by clustering the points within a predefined annulus around the candidate. The clustering approach is shown in Figure 2 as we adapt the filtering and clustering approaches to our purpose as described in section 6. In CellNet, the roads connecting the identified intersections are modeled using FastDTW, a fast and approximate implementation of dynamic time warping (DTW).
@@
Another approach to detect intersections involves initially identifying turning points by analyzing the difference in heading directions during individual trips, followed by clustering these turning points based on Euclidean distance ijgi4042446 (10). To reduce the incidence of falsely identified intersections, only clusters with multiple points are retained. Additionally, the heading directions into and out of each potential intersection are compared with each other to remove road bends. Similarly to CellNet, the tracks connecting the intersections are aligned using dynamic time warping.
@@
In COLTRANE (ConvolutiOnaL TRAjectory NEtwork) Prabowo_2019 (11), the authors propose a variation of the mean-shifting algorithm termed iterated trajectory mean shift (ITMS) to identify road centerlines. The road centerlines are treated as nodes, and the candidate edges are inferred based on the distance of the GPS points from the centerlines. Finally, the edges are classified and pruned with a convolutional neural network with custom input features derived from 2D histograms of location and directional velocities.
@@
The aforementioned methods have demonstrated strong performance in structured environments, such as cityscapes 10.1145/3234692 (9, 10, 11) and airport tarmacs Prabowo_2019 (11). However, the movement patterns of vehicles at construction sites differ from these environments, and include reversing and maneuvering into specific positions for loading and unloading of the materials, navigating through narrow roads and rugged terrain, wide open areas, variable speeds and frequently stopping and starting. To address these unique challenges, we propose a multi-step method that first identifies the intersections in the road network and subsequently connects the roads, thereby bridging the gap in current methodologies for construction site applications.
@@ 4 Data description
Median
Mean and std.
Range
Timestep
2 s
4.63 s $\pm$ 2.32 min
[0.88 s, 11.5 h]
Velocity (km/h)
8.33
9.36 $\pm$ 6.00
[0.0003, 25.93]
Distance (m)
17.13
34.97 $\pm$ 277.36
[0, 61241]
Nr points/trip
74
122.32 $\pm$ 188.90
[10, 3516]
@@
5 Data pre-processing
To enhance the quality of the data, the the following data pre-processing steps are performed:
@@ 5 Data pre-processing
      Velocity is zero: These are irrelevant for road inference as they do not represent movement.
@@
      The updates occur at the conclusion of trips (trip endpoints): These are often noisy because of two reasons: a) the vehicles move slowly right before loading/unloading, which creates a challenge in refining the location by the GPS receiver***GPS location is refined in the process called triangulation, which is based on the object movement., and b) the previous trip is considered ongoing until the truck is being loaded again, which can result in arbitrary movement around the loading site at the end of the trip.
@@
The hyperparameters and their default values are given in Table 3 in the Appendix.
@@ 6 Graph inference algorithm
[…]
@@ 6.2 Identification of candidate intersections
A cell $c_{i,j}$ is flagged as a preliminary intersection candidate if its corresponding $\Delta\phi_{i,j}$ exceeds a predefined threshold $\Delta\phi_{\mathrm{thr}}$. With the default cell resolution of $N_{\mathrm{res}}=5$ m, this typically results in multiple adjacent cells being flagged as potential intersections. To refine these results, the neighboring candidates are further merged with a simple clustering approach. Candidates within a distance of $D_{\mathrm{int\_clust}}=15$ m from each other are grouped together, and the central point of each group is marked as the final intersection candidate. Preliminary candidates that do not form a part of any cluster are disregarded under the assumption that the chosen $N_{\mathrm{res}}$ is sufficiently small for the intersections and their immediate surrounding to extend over several cells.
@@ 6.3 Validation of intersections
The next step consists of filtering out false positives among the candidate intersections using an adapted version of the method proposed in 10.1145/3234692 (9). This method discerns intersections from e.g., road bends by determining the number of roads leaving the candidate location for the intersection. In our version, we first subset the GPS points to those that lie within the circle bounded by the outer radius $R+L$ and apply DBSCAN clustering. The analysis is restricted to the GPS points that are directly connected to the intersection, i.e., they form a cluster with the points passing within a distance of $D_{\mathrm{passing}}=15$ m. Next, following 10.1145/3234692 (9), we define an extremity annulus bounded by two concentric circles with radii $R=30$ m and $R+L=55$ m around each candidate intersection (Fig. 2). These GPS points that lie within the annulus and are connected to intersection are clustered with a clustering distance of $D_{\mathrm{ext\_clust}}=30$ m. Each cluster with a size larger than $N_{\mathrm{ext\_clust}}=2$ represents a road passing through the extremity. The final set of intersections are those that have at least three such clusters in their extremity, representing three roads leaving the intersection.
Depending on the actual size of intersections, this step can be repeated with varying values of $R$ and $L$. To reduce computations for determining distances between the GPS points and intersections, we employed k-dimensional trees (k-d trees) that partition space into a hierarchical tree structure, significantly accelerating spatial queries. This approach allowed us to narrow down calculations only to those pairs of GPS points and intersection centers that are within a predefined maximum distance $R+L=55$ m from each other.
@@ 6.4 Identifying load and drop-off locations
Load and drop-off locations are also considered nodes. While load and drop-off locations are reported in the data for each individual trip, due to the manual nature of the logging process they are prone to uncertainty. To minimize the uncertainty, the reported locations are clustered as follows:
@@
    Loading points: The reported loading points are organized by the excavator ID and task ID. Within each group, the loading point is determined as the trimmed average of all positions. Loading points that are closer to each other than a predefined threshold, $D_{\mathrm{load/dump}}=100$ m, are merged.
@@
    Drop-off points: The reported drop-off points are clustered using density-based clustering (DBSCAN dbscan (12)) and the mean $(lat,lon)$ within each cluster is used to compute the drop-off location. The locations that are closer than $D_{\mathrm{load/dump}}=100$ m are merged.
@@ 6.5 Road inference
The process starts by identifying the trip sections that pass in proximity to any node, i.e., within a predefined radius of $D_{\mathrm{node}}=30$ m. For each of these sections, we mark the time step at which the GPS points are the closest to that node. The trips are then split into smaller segments at these time steps, resulting in segments that start at either the beginning of the trip or at a node, and similarly finish either at the end of the trip or at a node. The resulting segments are further grouped by their adjacent node pairs or singular nodes.
@@ 7 Results
The visual representation of the algorithm’s performance is presented in Fig. 3. Each of the four subplots show segments of the inferred graph, in the order from relatively straightforward to more complex scenarios. The parameters used are defined in the Appendix. The intersection validation step was run twice, with $R=30$ m and $R=100$ m (Fig. 4).
@@
In our further analysis we show the limitation of the algorithm in the presence of consistent noise. In Zone 3 four intersections are detected correctly (7, 8, 9, and 32). Number 33 is falsely identified as an intersection due to the noisy data, while an actual merging of the roads is located 100 m to the left. In this case, the noisy updates are more frequent than the correct outline of the road. Zone 4 presents the biggest challenge to the algorithm, with noisy data dominating several parts of the map, in the areas where the vehicles often change direction. In this case, the algorithm detects several false positive intersections, e.g., number 23, or 30.
@@ 9 Conclusions
Our findings highlight both the strengths and limitations of our proposed method. Notably, the algorithm demonstrates high level of proficiency in mapping large road networks and identifying road intersections and is able to achieve perfect precision and recall when the data quality is sufficient. On the other hand, the performance of our approach depends on the selection of parameters, indicating the necessity for a tailored approach rather than a one-size-fits-all solution across varied terrains. This automated graph construction approach will allow for improved coordination and management of units at construction site.
@@ Acknowledgements
This work is based upon the support from the Research Council of Norway under project Datadrevet Anleggsplass 309797 (English: Data-driven Construction Site). The data was collected by Skanska Norway AS and accessed through the Ditio apphttps://ditioapp.com/.
@@
    Katarzyna Michałowska: Conceptualization, methodology development (primary), coding and implementation, visualizations, analysis, writing (original draft),
@@
    Helga Margrete Bodahl Holmestad: Conceptualization, methodology development, coding and implementation, visualizations, writing (review and editing),
@@
    Signe Riemer-Sørensen: Conceptualization, methodology development, writing (review and editing), resources.
Data availability
The data used in this study are proprietary to Skanska AS and can be made available upon reasonable request.
@@
Authors state no conflict of interest.

### v3

@@ Abstract
We propose a new method for inferring roads from GPS trajectories to map construction sites. This task presents a unique challenge due to the erratic and non-standard movement patterns of construction machinery, which significantly diverge from typical vehicular traffic on established roads. Our proposed method first identifies intersections in the road network that serve as critical decision points, and then connects them with edges to produce a graph, which can subsequently be used for planning and task allocation. We demonstrate the approach by mapping roads at a real-life construction site in Norway. The method is validated on four increasingly complex segments of the map. In our tests, the method achieved perfect accuracy in detecting intersections and inferring roads in data with no or low noise, while its performance was reduced in areas with significant noise and consistently missing GPS updates.
@@ 1 Motivation
Reducing carbon emissions is a critical goal in mitigating the far-reaching impacts of climate change. In Norway, the building and construction sector contributes directly and indirectly to 15% of total greenhouse gas emissions (2019), with construction vehicles accounting for 1.5% of the total emissions (2021) BAE; Veikart.
In managing construction projects, it is crucial to effectively handle the complex interplay of resources, including personnel, materials, and machinery, to ensure their optimal interaction in terms of type, quantity, location, and time Dardouri2023. One viable strategy to address this issue is to maximize the use of heavy machinery, such as dump trucks and excavators, through optimal coordination and task allocation. According to Skanska Norge AS, the largest construction contractor in Norway, mass transporting units are idling 40–60% of the time during projects. This is a consequence of the conventional planning approach where excavators are considered more expensive than mass transportation units, and hence the latter are excessively deployed to avoid idling excavators, often leading to overcapacity and idling of transportation units instead. Improved coordination on construction sites, e.g., optimized real-time task allocation for dumpers, can reduce this overcapacity, leading to lower greenhouse gas emissions and saving both time and money. This is a critical improvement given the increasing size of construction projects involving hundreds of mass transporting units.
The first step in optimizing these operations is to acquire an up-to-date map of the construction site. It is imperative that the map generation be automated given the constantly evolving nature of the construction site and because manual mapping would be impractical and deprioritized, as emphasized in interviews with drivers, foremen, and construction site coordinators. Herein, we address this need by proposing an algorithm for road inference from Global Positioning System (GPS) trajectories recorded by vehicles on the construction site. Our algorithm generates a graph consisting of road intersections (nodes) and road centerlines (edges), which is a formulation that conveniently allows the subsequent use of standard resource planning and optimization techniques. Furthermore, utilizing movement data, as opposed to e.g., aerial images facilitates identification of temporary roads which are typically less visible or distinguishable in the terrain and does not require frequent collection of aerial images to capture the dynamic changes.
@@ 2 Related work
Automated road inference plays a pivotal role in various applications, such as urban planning, transportation management, and autonomous driving. The primary machine learning approaches for map generation are either from images, e.g., aerial photos, drone-, or satellite imagery Li2019; 8578594; Mattyus_2017_ICCV; BMVC2016_118; he2022td; SIEBERT20141, or from movement data, such as GPS trajectories 10.1145/3234692; ijgi4042446; Prabowo_2019; FU2017237; 10.1007/978-3-642-15300-6_5; he2018roadrunner; eftelioglu2022ring. The latter approach becomes particularly advantageous when paths are not easily distinguishable in images, e.g., blend into the surrounding environment, or when the tracks change more frequently than the aerial imagery is collected. Moreover, GPS data collection is cost-effective compared to aerial or satellite imagery, especially for repeated acquisitions over time, since it can leverage existing devices such as mobile phones.
@@
In the context of optimizing vehicle operations at construction sites, it is advantageous to model the road network as a graph. In such model, intersections are treated as key decision points, represented as nodes, while the roads themselves are depicted as the connecting edges between these nodes. Moreover, graph representation is more compact compared to alternatives such as raster images. We therefore direct our attention to works that detect road intersections and infer road centerlines, i.e., central paths along the roads.
@@
In CellNet 10.1145/3234692, the approach involves dividing the map into a grid and applying the mean-shifting algorithm on all GPS points within each cell, which results in identifying intersections, if they are present within the cell, or pinpointing the center of the stream of routes otherwise. To filter out the latter, it is verified that each potential intersection has at least three outgoing streams of routes by clustering the points within a predefined annulus around the candidate. The clustering approach is shown in Figure 2 as we adapt the filtering and clustering approaches to our purpose as described in section 6. In CellNet, the roads connecting the identified intersections are modeled using FastDTW, a fast and approximate implementation of dynamic time warping (DTW).
@@
Another approach to detect intersections involves initially identifying turning points by analyzing the difference in heading directions during individual trips, followed by clustering these turning points based on the Euclidean distance ijgi4042446. To reduce the incidence of falsely identified intersections, only clusters with multiple points are retained. Additionally, the heading directions into and out of each potential intersection are compared with each other to remove road bends. Similarly to CellNet, the tracks connecting the intersections are aligned using dynamic time warping.
@@
In COLTRANE (ConvolutiOnaL TRAjectory NEtwork) Prabowo_2019, the authors propose a variation of the mean-shifting algorithm termed iterated trajectory mean shift (ITMS) to identify road centerlines. The road centerlines are treated as nodes and the candidate edges are inferred based on the distance of the GPS points from the centerlines. Finally, the edges are classified and pruned with a convolutional neural network with custom input features derived from 2D histograms of location and directional velocities.
@@
The aforementioned methods have demonstrated strong performance in structured environments, such as cityscapes 10.1145/3234692; ijgi4042446; Prabowo_2019 and airport tarmacs Prabowo_2019. However, the movement patterns of vehicles at construction sites differ from these environments, and include reversing and maneuvering into specific positions for loading and unloading of the materials, navigating through narrow roads and rugged terrain, wide open areas, variable speeds and frequently stopping and starting. To address these unique challenges, we propose a multi-step method that first identifies the intersections in the road network and subsequently connects the roads, thereby bridging the gap in current methodologies for construction site applications.
@@ 4 Data description
Median Mean and std. Range
Timestep 2 s 4.63 s $\pm$ 2.32 min [0.88 s, 11.5 h]
Velocity (km/h) 8.33 9.36 $\pm$ 6.00 [0.0003, 25.93]
Distance (m) 17.13 34.97 $\pm$ 277.36 [0, 61241]
Nr points/trip 74 122.32 $\pm$ 188.90 [10, 3516]
@@
5 Data preprocessing
To enhance the quality of the data, the the following data preprocessing steps are performed:
@@ 5 Data preprocessing
      Velocity is zero: These are irrelevant for intersection identification and road inference as they do not represent movement. They are retained in the identification of the load and drop-off locations.
@@
      The updates occur at the conclusion of trips (trip endpoints): These are often noisy because of two reasons: a) the vehicles move slowly right before loading/unloading, which creates a challenge in refining the location by the GPS receiverGPS location is refined in the process called triangulation, which is based on the object movement., and b) the previous trip is considered ongoing until the truck is being loaded again, which can result in arbitrary movement around the loading site at the end of the trip.
@@
The hyperparameters and their default values are given in the Appendix.
@@ 6 Graph inference algorithm
In the first step, the map is partitioned into a grid (Fig. 1). Each cell is of size $N_{\mathrm{res}}\times N_{\mathrm{res}}=5\times 5$ m2, and is denoted as $c_{i,j}$, where $i$ and $j$ are the row and column indices, respectively. The GPS trajectories, divided into trips, can then be converted into movement directions $\Phi=\{\phi_{1},\phi_{2},...,\phi_{N}\}$ based on the latitude and longitude of consecutive GPS updates (similar to ijgi4042446). These directions are normalized to a $[0,\pi]$ range (00 to $180^{\circ}$), accounting for bi-directional vehicle movement and treating travel in either direction along the same path as equivalent. Finally, for each set $\Phi_{i,j}$ of points recorded within each cell $c_{i,j}$, the median $\tilde{\phi}_{i,j}$ is computed. This procedure essentially transforms the grid into a 2D histogram, where each cell represents the median of bi-directional vehicular movement.
@@ 6.2 Identification of candidate intersections
A cell $c_{i,j}$ is flagged as a preliminary intersection candidate if its corresponding $\Delta\phi_{i,j}$ exceeds a predefined threshold $\Delta\phi_{\mathrm{thr}}$. With the default cell resolution of $N_{\mathrm{res}}=5$ m, this typically results in multiple adjacent cells being flagged as potential intersections. To refine these results, the neighboring candidates are further merged with a simple clustering approach. Candidates within a distance of $D_{\mathrm{int\_clust}}=15$ m from each other are grouped together, and the central point of each group is marked as the final intersection candidate. Preliminary candidates that do not form a part of any cluster are excluded under the assumption that the chosen $N_{\mathrm{res}}$ is small enough for the intersections and their immediate surroundings to extend over several cells.
@@ 6.3 Validation of intersections
The next step consists of filtering out false positives among the candidate intersections using an adapted version of the method proposed in 10.1145/3234692. This method discerns intersections from e.g., road bends by determining the number of roads leaving the candidate location for the intersection. In our version, we first subset the GPS points to those that lie within the circle bounded by the outer radius $R+L$ and apply DBSCAN clustering. The analysis is restricted to the GPS points that are directly connected to the intersection, i.e., they form a cluster with the points passing within a distance of $D_{\mathrm{passing}}=15$ m. Next, following 10.1145/3234692, we define an extremity annulus bounded by two concentric circles with radii $R=30$ m and $R+L=55$ m around each candidate intersection (Fig. 2). These GPS points located within the annulus and are connected to the intersection are clustered with a clustering distance of $D_{\mathrm{ext\_clust}}=30$ m. Each cluster of a size larger than $N_{\mathrm{ext\_clust}}=2$ represents a road that passes through the extremity. The final set of intersections are those that have at least three such clusters in their extremity, representing three roads leaving the intersection.
Depending on the actual size of intersections, this step can be repeated with varying values of $R$ and $L$. To reduce computations for determining distances between the GPS points and intersections, we employed k-dimensional trees (k-d trees) that partition space into a hierarchical tree structure, significantly accelerating spatial queries. This approach allowed us to narrow down calculations to only those pairs of GPS points and intersection centers that are within a predefined maximum distance $R+L=55$ m from each other.
@@ 6.4 Identifying load and drop-off locations
Load and drop-off locations are also considered nodes. Although load and drop-off locations are reported in the data for each individual trip, due to the manual nature of the logging process, they are prone to uncertainty. To minimize the uncertainty, the reported locations are clustered as follows:
@@
    Loading points: The reported loading points are organized by excavator ID and task ID. Within each group, the loading point is determined as the trimmed average of all positions. The loading points that are closer to each other than a predefined threshold, $D_{\mathrm{load/dump}}=100$ m, are merged.
@@
    Drop-off points: The reported drop-off points are clustered using density-based clustering (DBSCAN dbscan) and the mean $(lat,lon)$ within each cluster is used to compute the drop-off location. The locations that are closer than $D_{\mathrm{load/dump}}=100$ m are merged.
@@ 6.5 Road inference
The process starts by identifying the sections of the trip that pass in proximity to any node, i.e., within a predefined radius of $D_{\mathrm{node}}=30$ m. For each of these sections, we mark the time step at which the GPS points are the closest to that node. The trips are then split into smaller segments at these time steps, resulting in segments that start at either the beginning of the trip or at a node, and similarly finish either at the end of the trip or at a node. The resulting segments are further grouped by their adjacent node pairs or singular nodes.
@@ 7 Results
The visual representation of the algorithm’s performance is presented in Fig. 3. Each of the four subplots shows segments of the inferred graph in the order from relatively straightforward to more complex scenarios. The parameters used are defined in the Appendix. The intersection validation step was run twice, with $R=30$ m and $R=100$ m (Fig. 4).
@@
In our further analysis, we show the limitation of the algorithm in the presence of consistent noise. In Zone 3 four intersections are detected correctly (7, 8, 9, and 32). Number 33 is falsely identified as an intersection due to the noisy data, while an actual merging of the roads is located 100 m to the left. In this case, the noisy updates are more frequent than the correct outline of the road. Zone 4 presents the biggest challenge to the algorithm, with noisy data dominating several parts of the map, in the areas where the vehicles often change direction. In this case, the algorithm detects several false positive intersections, e.g., number 23, or 30.
@@ 9 Conclusions
We have introduced a novel method for road inference from GPS data, which constructs a detailed graph of road networks with nodes representing intersections, load points, and drop-off points, while edges correspond to the connecting roads. The algorithm demonstrates a high level of proficiency in mapping large-scale road networks and identifying road intersections and achieves perfect precision and recall when data quality is high. However, it can be sensitive to noise, particularly for irregular GPS updates, e.g., when the signal is lost as a vehicle passes through a tunnel. Our approach incorporates a selection of physically meaningful and understandable parameters that offer flexibility for tuning across diverse terrains. This automated graph construction approach will allow for improved coordination and management of units at construction sites.
@@ Acknowledgements
This work is based upon the support from the Research Council of Norway under the Datadrevet Anleggsplass 309797 (English: Data-driven Construction Site). Data were collected by Skanska Norway AS and accessed through the Ditio app (https://ditioapp.com/).
@@
    Katarzyna Michałowska: Conceptualization, methodology, data curation, software, investigation, formal analysis, validation, visualization, writing (original draft),
@@
    Helga Margrete Bodahl Holmestad: Conceptualization, methodology, data curation, software, investigation, visualization, writing (review and editing),
@@
    Signe Riemer-Sørensen: Conceptualization, methodology, writing (review and editing), resources, project administration, funding acquisition.
Data and code availability
The data used in this study are proprietary to Skanska AS and can be made available upon reasonable request. The code developed in this study is available in this GitHub repository: https://github.com/katarzynamichalowska/road-graph-generator.
@@
The authors declare no conflict of interest.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT to improve the clarity of the language in certain sections of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
[…]